OpenStudy (anonymous):

A lawn mowers tank is 2/3 full. it has 1/2 gallon of gas in it. How many gallons does it hold?

OpenStudy (bibby):

\(\huge \frac{2}{3}tankcapacity=\frac{1}{2} \) set up an equation

OpenStudy (anonymous):

So would i Divide 2/3 divided by 1/2? @bibby

OpenStudy (bibby):

no

OpenStudy (bibby):

you want to find the tank capacity so that's our X. now we know that 2/3 OF (of=multiply) x is 1/2 now we can divide both sides by 2/3

OpenStudy (anonymous):

Ohhh ok thanks